# Towards Principled Superhuman AI for Multiplayer Symmetric Games

### v1

@@ A.1 $3$-player majority and minority game
for some absolute constant $C>0$. Thus we finish the proof. ∎
Appendix C Proofs for Section 5.2
C.1 Proof of Fact 5.2
Proof of Fact 5.2.
The proof of Fact 5.2 follows directly from the construction in Theorem 5.6 with $n=3$, $\mathcal{A}=\{0,1\}$ and $\Delta_{T}=1$. ∎
C.2 Proof of Theorem 5.3
@@
$\quad+\underbrace{\sum_{t\in\mathcal{I}}\left(U_{1}(a_{0},a^{t}_{-1})-U_{1}(x^{t},a^{t}_{-1})\right)}_{\rm(ii)}+\sum_{t\in\mathcal{I}}\left(U_{1}(x^{t},a^{t}_{-1})-u^{t}(x^{t})\right)$
@@
where the last inequality follows from Theorem 1 in (Daniely et al.,, 2015).
@@
$\leq 2V_{\mathcal{I}}|\mathcal{I}|+\sum_{t\in\mathcal{I}}\left(u^{t}(a_{0})-U_{1}(a_{0},a^{t}_{-1})\right)+C(\sqrt{\log A}+\log T)\sqrt{|\mathcal{I}|}+\sum_{t\in\mathcal{I}}\left(U_{1}(x^{t},a^{t}_{-1})-u^{t}(x^{t})\right).$
@@
$\leq 2V_{\mathcal{I}_{j}}|\mathcal{I}|+\sum_{t\in\mathcal{I}_{j}}\left(u^{t}(a_{0})-U_{1}(a_{0},a^{t}_{-1})\right)+C(\sqrt{\log A}+\log T)\sqrt{|\mathcal{I}|}+\sum_{t\in\mathcal{I}_{j}}\left(U_{1}(x^{t},a^{t}_{-1})-u^{t}(x^{t})\right).$
@@
$\leq 2V_{T}|\mathcal{I}|+\underbrace{\sum^{T}_{t=1}\left(u^{t}(a_{0})-U_{1}(a_{0},a^{t}_{-1})\right)}_{\rm(iii)}+C(T/\sqrt{|\mathcal{I}|})\cdot(\sqrt{\log A}+\log T)+\underbrace{\sum^{T}_{t=1}\left(U_{1}(x^{t},a^{t}_{-1})-u^{t}(x^{t})\right)}_{\rm(iv)}.$
@@
C.3 Proof of Theorem 5.4
@@
Thus we finish the proof. ∎
Appendix D Fundamental limits
Upon examining Theorem 5.3 alongside Theorem 5.4, it becomes apparent that Theorem 5.3 benchmarks against a more stringent standard (i.e., the dynamic oracle) and incurs a larger regret of $V^{1/3}_{T}T^{2/3}$, while Theorem 5.4 sets its comparison against a baseline metric (i.e., the average payoff) and attains a smaller regret of $V_{T}$. Regarding this observation, one might aspire to devise an algorithm whose payoff satisfies: $\sum^{T}_{t=1}u^{t}(x^{t})\geq\sum_{t=1}^{T}\max_{a\in\mathcal{A}}u^{t}(a)-\tilde{O}(V_{T}).$ However, Theorem 5.5 and Theorem 5.6 demonstrate that such a goal is unattainable, by exploring the fundamental limits faced when competing against non-stationary opponents.
Theorem 5.5 shows, when contending with non-stationary opponent, the optimal algorithm must incur a dynamic regret at least order of $V^{1/3}_{T}T^{2/3}$, closing off the possibility of attaining a better $V_{T}$ rate. It’s noteworthy that a similar lower bound for dynamic regret has already been established under broader conditions Besbes et al., (2014). The distinction of Theorem 5.5 lies in further restricting the hard problems to be symmetric games, implying that the structure of symmetric game does not offer an advantage in improving dynamic regret in the worst case. By comparing this lower bound with Theorem 5.3, it is evident that $\mathrm{SAOL}^{\mathcal{H}}$ is demonstrated to be minimax optimal, albeit with the inclusion of some logarithmic factors.
@@ Appendix D Fundamental limits
$U^{(3)}_{1}(a,b,c):=\begin{cases}\text{payoff for $3$-player majority game}&\text{if }a,b,c\in\{0,1\}\\ -2&\text{if }a\notin\{0,1\},b,c\in\{0,1\}\\ \text{defined by symmetric}&\text{o.w.}\end{cases}$
@@
Lemma D.1.
Let $f:\{-2,0,1\}^{|\mathcal{T}_{j}|\times A}\rightarrow[0,M]$ be any bounded real function defined on the payoff matrices $R$. Then, for any $\delta_{j}\in\{1,2\}$, ${\epsilon}\leq 1/4$:
@@
By Lemma D.1 with $f=N^{j}_{-\delta_{j}+2}$, we have
@@
$\mathbb{E}^{j}_{\delta_{j}}[u^{t}(x^{t})]$ $=-2\mathbb{P}^{j}_{\delta_{j}}(x^{t}\notin\{0,1\})+(-2{\epsilon}-4{\epsilon}^{2})\mathbb{P}^{j}_{\delta_{j}}(x^{t}=\delta_{j}-1)+(2{\epsilon}-4{\epsilon}^{2})\mathbb{P}^{j}_{\delta_{j}}(x^{t}=-\delta_{j}+2)$
$\leq(-2{\epsilon}-4{\epsilon}^{2})\mathbb{P}^{j}_{\delta_{j}}(x^{t}\neq-\delta_{j}+2)+(2{\epsilon}-4{\epsilon}^{2})\mathbb{P}^{j}_{\delta_{j}}(x^{t}=-\delta_{j}+2)$
$=-2{\epsilon}-4{\epsilon}^{2}+4{\epsilon}\cdot\mathbb{P}^{j}_{\delta_{j}}(x^{t}=-\delta_{j}+2),$
@@
$\mathbb{E}^{j}_{\delta_{j}}\left[\sum_{t\in\mathcal{T}_{j}}u^{t}(x^{t})\right]$ $\leq(-2{\epsilon}-4{\epsilon}^{2})|\mathcal{T}_{j}|+4{\epsilon}\cdot\mathbb{E}^{j}_{\delta_{j}}[N^{j}_{-\delta_{j}+2}]$
$\leq(-2{\epsilon}-4{\epsilon}^{2})|\mathcal{T}_{j}|+4{\epsilon}\cdot\mathbb{E}^{j}_{0}[N^{j}_{-\delta_{j}+2}]+8{\epsilon}^{2}|\mathcal{T}_{j}|\sqrt{\Delta_{T}}.$ (by (3))
@@
$\frac{1}{2}\mathbb{E}^{j}_{1}\left[\sum_{t\in\mathcal{T}_{j}}u^{t}(x^{t})\right]+\frac{1}{2}\mathbb{E}^{j}_{2}\left[\sum_{t\in\mathcal{T}_{j}}u^{t}(x^{t})\right]\leq(-2{\epsilon}-4{\epsilon}^{2})|\mathcal{T}_{j}|+2{\epsilon}|\mathcal{T}_{j}|+8{\epsilon}^{2}|\mathcal{T}_{j}|\sqrt{\Delta_{T}}$ (4)
@@
$\leq\sum^{m}_{j=1}((-2{\epsilon}-4{\epsilon}^{2})|\mathcal{T}_{j}|+2{\epsilon}|\mathcal{T}_{j}|+8{\epsilon}^{2}|\mathcal{T}_{j}|\sqrt{\Delta_{T}})$
$=-4{\epsilon}^{2}T+8{\epsilon}^{2}T\sqrt{\Delta_{T}}.$
@@
$\mathbb{E}_{{\text{Alg}}}[{\text{D-Reg}(T)}]$ $=(2{\epsilon}-4{\epsilon}^{2})T-\mathbb{E}_{{\text{Alg}}}\left[\sum^{T}_{t=1}u^{t}(x^{t})\right]$
$\geq(2{\epsilon}-4{\epsilon}^{2})T-(-4{\epsilon}^{2}T+8{\epsilon}^{2}T\sqrt{\Delta_{T}})$
$=2{\epsilon}T-8{\epsilon}^{2}T\sqrt{\Delta_{T}}$
$=2{\epsilon}T(1-4{\epsilon}\sqrt{\Delta_{T}})$
$\geq{\epsilon}T$
$=\min\left\{\frac{1}{8\sqrt{\Delta_{T}}},\frac{V_{T}\Delta_{T}}{T}\right\}T.$
@@
$\mathbb{E}_{{\text{Alg}}}[{\text{D-Reg}(T)}]\geq CV_{T}^{1/3}T^{2/3},$
which completes the proof. ∎
We prove Lemma D.1 in the following.
Proof of Lemma D.1.
@@
$\mathbb{P}_{0}(R_{t}=[-2,0,-2,\ldots,-2]\mid R^{t-1})=\mathbb{P}_{0}(R_{t}=[0,-2,-2,\ldots,-2]\mid R^{t-1})=1/4$
$\mathbb{P}_{0}(R_{t}=[1,1,-2,\ldots,-2]\mid R^{t-1})=1/2.$
@@
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[-2,0,-2,\ldots,-2]\mid R^{t-1})=(1/2+{\epsilon})^{2}$
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[0,-2,-2,\ldots,-2]\mid R^{t-1})=(1/2-{\epsilon})^{2}$
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[1,1,-2,\ldots,-2]\mid R^{t-1})=2(1/2+{\epsilon})(1/2-{\epsilon}).$
@@
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[-2,0,-2,\ldots,-2]\mid R^{t-1})=(1/2-{\epsilon})^{2}$
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[0,-2,-2,\ldots,-2]\mid R^{t-1})=(1/2+{\epsilon})^{2}$
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[1,1,-2,\ldots,-2]\mid R^{t-1})=2(1/2+{\epsilon})(1/2-{\epsilon}).$
@@
In Case 1, we have $u^{t}(0)=0$ and $u^{t}(a)=-2$ for all $a\neq 0$. In Case 2, we have $u^{t}(1)=0$ and $u^{t}(a)=-2$ for all $a\neq 1$. At the beginning of each batch, one of these scenarios is randomly selected (with equal probability) and remains constant throughout that batch.
@@
$\frac{1}{2}\mathbb{E}^{j}_{1}\left[u^{t^{j,1}}(x^{t^{j,1}})\right]+\frac{1}{2}\mathbb{E}^{j}_{2}\left[u^{t^{j,1}}(x^{t^{j,1}})\right]$ $=-\left(\mathbb{P}^{j}_{1}(x^{t^{j,1}}\neq 0)+\mathbb{P}^{j}_{2}(x^{t^{j,1}}\neq 1)\right)$
$=-\left(\mathbb{P}(x^{t^{j,1}}\neq 0)+\mathbb{P}(x^{t^{j,1}}\neq 1)\right)\leq-1,$
@@
$\mathbb{E}_{{\text{Alg}}}\left[\sum^{m}_{j=1}\sum_{t\in\mathcal{T}_{j}}u^{t}(x^{t})\right]\leq-m=-T/\Delta_{T}.$
Choosing $\Delta_{T}=2T/V_{T}$, we have
@@
Appendix E Experiments Details
E.1 Algorithms
We refer readers to Algorithm 2-5 for detailed implementation of algorithms in the experiment. For MV, we choose $\eta=1$. For SDG, we choose $\eta=2$.
@@ E.1 Algorithms
0:  Number of iterations $T$, action space $\mathcal{A}$, learning rate $\eta_{t}=\eta\sqrt{\frac{\log{|\mathcal{A}|}}{t}}$, number of players $n$, and initialize policy $\pi_{0}$.
@@
2:     Sample actions $a^{t-1}_{i}\sim\pi_{t-1}$ for $i=2,\ldots,n$. Denote $a^{t-1}_{-1}:=(a^{t-1}_{2},\ldots,a^{t-1}_{n})$.
@@
$\pi_{t}(a)\propto\pi_{t-1}(a){\sf exp}{\{\eta_{t}U_{1}(a,a^{t-1}_{-1})\}},\forall a\in\mathcal{A}.$
@@
0:  Number of iterations $T$, action space $\mathcal{A}$, learning rate $\eta_{t}=\eta\sqrt{\frac{\log{|\mathcal{A}|}}{t}}$, number of players $n$, initialize policy $\pi_{0}$, human policy $\pi_{{\text{human}}}$ and regularization parameter $\lambda$.
@@
2:     Sample actions $a^{t-1}_{i}\sim\pi_{t-1}$ for $i=2,\ldots,n$. Denote $a^{t-1}_{-1}:=(a^{t-1}_{2},\ldots,a^{t-1}_{n})$.
@@
$\pi_{t}(a)\propto{\sf exp}{\left\{\frac{\log\pi_{0}(a)+\sum_{\tau<t}\eta_{\tau}U_{1}(a,a^{\tau}_{-1})+\lambda\sum_{\tau<t}\eta_{\tau}\log\pi_{{\text{human}}}(a)}{1+\lambda\sum_{\tau<t}\eta_{\tau}}\right\}},\forall a\in\mathcal{A}.$
@@
Algorithm 4 Best Response to $\pi_{{\text{human}}}$
0:  Number of iterations $T$, action space $\mathcal{A}$, learning rate $\eta_{t}=\eta\sqrt{\frac{\log{|\mathcal{A}|}}{t}}$, number of players $n$, and initialize policy $\pi_{0}$.
@@
2:     Sample actions $a^{t-1}_{i}\sim\pi_{{\text{human}}}$ for $i=2,\ldots,n$. Denote $a^{t-1}_{-1}:=(a^{t-1}_{2},\ldots,a^{t-1}_{n})$.
@@
$\pi_{t}(a)\propto\pi_{t-1}(a){\sf exp}{\{\eta_{t}U_{1}(a,a^{t-1}_{-1})\}},\forall a\in\mathcal{A}.$
@@
Algorithm 5 Exploiter for $\pi$
0:  Number of iterations $T$, action space $\mathcal{A}$, learning rate $\eta_{t}=\eta\sqrt{\frac{\log{|\mathcal{A}|}}{t}}$, number of players $n$, and initialize policy $\pi_{0}$.
@@
2:     Sample actions $a^{t-1}_{1}\sim\pi$.
3:     Sample actions $a^{t-1}_{i}\sim\pi_{t-1}$ for $i=2,\ldots,n$. Denote $a^{t-1}_{-1}:=(a^{t-1}_{2},\ldots,a^{t-1}_{n})$.
@@
$\pi_{t}(a)\propto\pi_{t-1}(a){\sf exp}{\left\{-\frac{\eta_{t}}{n-1}\sum^{n}_{i=2}U_{1}\left(a^{t-1}_{1},a^{t-1}_{-1}[:i-1],a,a^{t-1}_{-1}[i+1:]\right)\right\}},\forall a\in\mathcal{A}.$

### v2

@@ C.1 Guarantees for efficient algorithms
for some absolute constant $C>0$. Thus we finish the proofs. ∎
C.1.2 Proof of Theorem 5.2
In this section, we establish guarantees for $\mathrm{SAOL}^{\mathcal{H}}$ when facing adaptive opponents.
@@
$\quad+\underbrace{\sum_{t\in\mathcal{I}}\left(U_{1}(a_{0},a^{t}_{-1})-U_{1}(x^{t},a^{t}_{-1})\right)}_{\rm(ii)}+\sum_{t\in\mathcal{I}}\left(U_{1}(x^{t},a^{t}_{-1})-u^{t}(x^{t})\right).$
@@
where the last inequality follows from Theorem 1 in Daniely et al. (2015).
@@
$\leq 2V_{\mathcal{I}}|\mathcal{I}|+\sum_{t\in\mathcal{I}}\left(u^{t}(a_{0})-U_{1}(a_{0},a^{t}_{-1})\right)+C(\sqrt{\log A}+\log T)\sqrt{|\mathcal{I}|}$
$\quad+\sum_{t\in\mathcal{I}}\left(U_{1}(x^{t},a^{t}_{-1})-u^{t}(x^{t})\right).$
@@
$\leq 2V_{\mathcal{I}_{j}}|\mathcal{I}|+\sum_{t\in\mathcal{I}_{j}}\left(u^{t}(a_{0})-U_{1}(a_{0},a^{t}_{-1})\right)+C(\sqrt{\log A}+\log T)\sqrt{|\mathcal{I}|}$
$\quad+\sum_{t\in\mathcal{I}_{j}}\left(U_{1}(x^{t},a^{t}_{-1})-u^{t}(x^{t})\right).$
@@
$\leq 2V_{T}|\mathcal{I}|+\underbrace{\sum^{T}_{t=1}\left(u^{t}(a_{0})-U_{1}(a_{0},a^{t}_{-1})\right)}_{\rm(iii)}+C(T/\sqrt{|\mathcal{I}|})\cdot(\sqrt{\log A}+\log T)$
$\quad+\underbrace{\sum^{T}_{t=1}\left(U_{1}(x^{t},a^{t}_{-1})-u^{t}(x^{t})\right)}_{\rm(iv)}.$
@@
Finally, by the definition of ${u_{\text{avg}}(T)}$ and $u^{\dagger}$, we show that
${u_{\text{avg}}(T)}\geq u^{\dagger}-CV^{1/3}_{T}T^{-1/3}\left(\sqrt{\log(A/\delta)}+\log T\right)$
for some absolute constant $C$.
C.1.3 Proof of Theorem 5.3
In this section, we establish guarantees for behavior cloning when facing adaptive opponents.
@@
Finally, by the definition of ${u_{\text{avg}}(T)}$, we finish the proofs. ∎
C.2 Matching lower bounds
Upon examining Theorem 5.2 alongside Theorem 5.3, it becomes apparent that Theorem 5.2 benchmarks against a more stringent standard (i.e., the dynamic oracle) and incurs a larger error of $V^{1/3}_{T}T^{-1/3}$, while Theorem 5.3 sets its comparison against a baseline metric (i.e., the average payoff) and attains a smaller error of $V_{T}/T$. Regarding this observation, one might aspire to devise an algorithm whose payoff satisfies: ${u_{\text{avg}}(T)}\geq u^{\dagger}-\tilde{O}(V_{T}/T).$ However, Theorem 5.4 and Theorem 5.5 demonstrate that such a goal is unattainable, by exploring the fundamental limits faced when competing against non-stationary opponents.
Theorem 5.4 shows, when contending with non-stationary opponent, the optimal algorithm must incur a dynamic regret at least order of $V^{1/3}_{T}T^{2/3}$, closing off the possibility of attaining a better $V_{T}$ rate. It’s noteworthy that a similar lower bound for dynamic regret has already been established under broader conditions Besbes et al. (2014). The distinction of Theorem 5.4 lies in further restricting the hard problems to be symmetric games, implying that the structure of symmetric game does not offer an advantage in improving dynamic regret in the worst case. By comparing this lower bound with Theorem 5.2, it is evident that $\mathrm{SAOL}^{\mathcal{H}}$ is demonstrated to be minimax optimal, albeit with the inclusion of some logarithmic factors.
@@ C.2 Matching lower bounds
$U^{(3)}_{1}(a,b,c):=\begin{cases}\text{payoff for $3$-player majority game}&\text{if }a,b,c\in\{0,1\}\\ -1&\text{if }a\notin\{0,1\},b,c\in\{0,1\}\\ \text{defined by symmetric}&\text{o.w.}\end{cases}$
@@
Lemma C.1.
Let $f:\{-1,0,1/2\}^{|\mathcal{T}_{j}|\times A}\rightarrow[0,M]$ be any bounded real function defined on the payoff matrices $R$. Then, for any $\delta_{j}\in\{1,2\}$, ${\epsilon}\leq 1/4$:
@@
By Lemma C.1 with $f=N^{j}_{-\delta_{j}+2}$, we have
@@
$\mathbb{E}^{j}_{\delta_{j}}[u^{t}(x^{t})]$ $=-\mathbb{P}^{j}_{\delta_{j}}(x^{t}\notin\{0,1\})+(-{\epsilon}-2{\epsilon}^{2})\mathbb{P}^{j}_{\delta_{j}}(x^{t}=\delta_{j}-1)+({\epsilon}-2{\epsilon}^{2})\mathbb{P}^{j}_{\delta_{j}}(x^{t}=-\delta_{j}+2)$
$\leq(-{\epsilon}-2{\epsilon}^{2})\mathbb{P}^{j}_{\delta_{j}}(x^{t}\neq-\delta_{j}+2)+({\epsilon}-2{\epsilon}^{2})\mathbb{P}^{j}_{\delta_{j}}(x^{t}=-\delta_{j}+2)$
$=-{\epsilon}-2{\epsilon}^{2}+2{\epsilon}\cdot\mathbb{P}^{j}_{\delta_{j}}(x^{t}=-\delta_{j}+2),$
@@
$\mathbb{E}^{j}_{\delta_{j}}\left[\sum_{t\in\mathcal{T}_{j}}u^{t}(x^{t})\right]$ $\leq(-{\epsilon}-2{\epsilon}^{2})|\mathcal{T}_{j}|+2{\epsilon}\cdot\mathbb{E}^{j}_{\delta_{j}}[N^{j}_{-\delta_{j}+2}]$
$\leq(-{\epsilon}-2{\epsilon}^{2})|\mathcal{T}_{j}|+2{\epsilon}\cdot\mathbb{E}^{j}_{0}[N^{j}_{-\delta_{j}+2}]+4{\epsilon}^{2}|\mathcal{T}_{j}|\sqrt{\Delta_{T}}.$ (by (4))
@@
$\frac{1}{2}\mathbb{E}^{j}_{1}\left[\sum_{t\in\mathcal{T}_{j}}u^{t}(x^{t})\right]+\frac{1}{2}\mathbb{E}^{j}_{2}\left[\sum_{t\in\mathcal{T}_{j}}u^{t}(x^{t})\right]\leq(-{\epsilon}-2{\epsilon}^{2})|\mathcal{T}_{j}|+{\epsilon}|\mathcal{T}_{j}|+4{\epsilon}^{2}|\mathcal{T}_{j}|\sqrt{\Delta_{T}}.$ (5)
@@
$\leq\sum^{m}_{j=1}((-{\epsilon}-2{\epsilon}^{2})|\mathcal{T}_{j}|+{\epsilon}|\mathcal{T}_{j}|+4{\epsilon}^{2}|\mathcal{T}_{j}|\sqrt{\Delta_{T}})$
$=-2{\epsilon}^{2}T+4{\epsilon}^{2}T\sqrt{\Delta_{T}}.$
@@
$\mathbb{E}_{{\text{Alg}}}[{\text{D-Reg}(T)}]$ $=({\epsilon}-2{\epsilon}^{2})T-\mathbb{E}_{{\text{Alg}}}\left[\sum^{T}_{t=1}u^{t}(x^{t})\right]$
$\geq({\epsilon}-2{\epsilon}^{2})T-(-2{\epsilon}^{2}T+4{\epsilon}^{2}T\sqrt{\Delta_{T}})$
$={\epsilon}T-4{\epsilon}^{2}T\sqrt{\Delta_{T}}$
$={\epsilon}T(1-4{\epsilon}\sqrt{\Delta_{T}})$
$\geq\frac{1}{2}{\epsilon}T$
$=\frac{1}{2}\min\left\{\frac{1}{8\sqrt{\Delta_{T}}},\frac{V_{T}\Delta_{T}}{T}\right\}T.$
@@
$\mathbb{E}_{{\text{Alg}}}[{\text{D-Reg}(T)}]\geq CV_{T}^{1/3}T^{2/3}.$
Recall the definition of ${u_{\text{avg}}(T)}$ and $u^{\dagger}$, we then finish the proofs. ∎
We prove Lemma C.1 in the following.
Proof of Lemma C.1.
@@
$\mathbb{P}_{0}(R_{t}=[-1,0,-1,\ldots,-1]\mid R^{t-1})=\mathbb{P}_{0}(R_{t}=[0,-1,-1,\ldots,-1]\mid R^{t-1})=1/4$
$\mathbb{P}_{0}(R_{t}=[1/2,1/2,-1,\ldots,-1]\mid R^{t-1})=1/2.$
@@
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[-1,0,-1,\ldots,-1]\mid R^{t-1})=(1/2+{\epsilon})^{2}$
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[0,-1,-1,\ldots,-1]\mid R^{t-1})=(1/2-{\epsilon})^{2}$
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[1/2,1/2,-1,\ldots,-1]\mid R^{t-1})=2(1/2+{\epsilon})(1/2-{\epsilon}).$
@@
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[-1,0,-1,\ldots,-1]\mid R^{t-1})=(1/2-{\epsilon})^{2}$
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[0,-1,-1,\ldots,-1]\mid R^{t-1})=(1/2+{\epsilon})^{2}$
$\mathbb{P}^{j}_{\delta_{j}}(R_{t}=[1/2,1/2,-1,\ldots,-1]\mid R^{t-1})=2(1/2+{\epsilon})(1/2-{\epsilon}).$
@@
In Case 1, we have $u^{t}(0)=0$ and $u^{t}(a)=-1$ for all $a\neq 0$. In Case 2, we have $u^{t}(1)=0$ and $u^{t}(a)=-1$ for all $a\neq 1$. At the beginning of each batch, one of these scenarios is randomly selected (with equal probability) and remains constant throughout that batch.
@@
$\frac{1}{2}\mathbb{E}^{j}_{1}\left[u^{t^{j,1}}(x^{t^{j,1}})\right]+\frac{1}{2}\mathbb{E}^{j}_{2}\left[u^{t^{j,1}}(x^{t^{j,1}})\right]$ $=-\frac{1}{2}\left(\mathbb{P}^{j}_{1}(x^{t^{j,1}}\neq 0)+\mathbb{P}^{j}_{2}(x^{t^{j,1}}\neq 1)\right)$
$=-\frac{1}{2}\left(\mathbb{P}(x^{t^{j,1}}\neq 0)+\mathbb{P}(x^{t^{j,1}}\neq 1)\right)\leq-\frac{1}{2},$
@@
$\mathbb{E}_{{\text{Alg}}}\left[\sum^{m}_{j=1}\sum_{t\in\mathcal{T}_{j}}u^{t}(x^{t})\right]\leq-\frac{m}{2}=-\frac{T}{2\Delta_{T}}.$
Choosing $\Delta_{T}=T/V_{T}$, we have
@@
Recall the definition of ${u_{\text{avg}}(T)}$, we then finish the proofs. ∎
Appendix D Experiments Details
In this section, we provide additional details for our experiments.
D.1 Algorithms
We refer readers to Algorithm 3-6 for detailed implementation of algorithms in the experiment. For MV, we choose $\eta=1$. For SDG, we choose $\eta=2$.
@@ D.1 Algorithms
0:  Number of iterations $T$, action space $\mathcal{A}$, learning rate $\eta_{t}=\eta\sqrt{\frac{\log{|\mathcal{A}|}}{t}}$, number of players $n$, and initialize strategy $x_{0}$.
@@
2:     Sample actions $a^{t-1}_{i}\sim x_{t-1}$ for $i=2,\ldots,n$. Denote $a^{t-1}_{-1}:=(a^{t-1}_{2},\ldots,a^{t-1}_{n})$.
@@
$x_{t}(a)\propto x_{t-1}(a){\sf exp}{\{\eta_{t}U_{1}(a,a^{t-1}_{-1})\}},\forall a\in\mathcal{A}.$
@@
0:  Number of iterations $T$, action space $\mathcal{A}$, learning rate $\eta_{t}=\eta\sqrt{\frac{\log{|\mathcal{A}|}}{t}}$, number of players $n$, initialize strategy $x_{0}$, meta-strategy $y_{\text{meta}}$, and regularization parameter $\lambda$.
@@
2:     Sample actions $a^{t-1}_{i}\sim x_{t-1}$ for $i=2,\ldots,n$. Denote $a^{t-1}_{-1}:=(a^{t-1}_{2},\ldots,a^{t-1}_{n})$.
@@
$x_{t}(a)\propto{\sf exp}{\left\{\frac{\log x_{0}(a)+\sum_{\tau<t}\eta_{\tau}U_{1}(a,a^{\tau}_{-1})+\lambda\sum_{\tau<t}\eta_{\tau}\log y_{\text{meta}}(a)}{1+\lambda\sum_{\tau<t}\eta_{\tau}}\right\}},\forall a\in\mathcal{A}.$
@@
Algorithm 5 Hedge
0:  Number of iterations $T$, action space $\mathcal{A}$, learning rate $\eta_{t}=\eta\sqrt{\frac{\log{|\mathcal{A}|}}{t}}$, number of players $n$, and initialize strategy $x_{0}$.
@@
2:     Sample actions $a^{t-1}_{i}\sim y_{\text{meta}}$ for $i=2,\ldots,n$. Denote $a^{t-1}_{-1}:=(a^{t-1}_{2},\ldots,a^{t-1}_{n})$.
@@
$x_{t}(a)\propto x_{t-1}(a){\sf exp}{\{\eta_{t}U_{1}(a,a^{t-1}_{-1})\}},\forall a\in\mathcal{A}.$
@@
Algorithm 6 Exploiter for strategy $x$
0:  Number of iterations $T$, action space $\mathcal{A}$, learning rate $\eta_{t}=\eta\sqrt{\frac{\log{|\mathcal{A}|}}{t}}$, number of players $n$, and initialize strategy $x_{0}$.
@@
2:     Sample actions $a^{t-1}_{1}\sim x$.
3:     Sample actions $a^{t-1}_{i}\sim x_{t-1}$ for $i=2,\ldots,n$. Denote $a^{t-1}_{-1}:=(a^{t-1}_{2},\ldots,a^{t-1}_{n})$.
@@
$x_{t}(a)\propto x_{t-1}(a){\sf exp}{\left\{-\frac{\eta_{t}}{n-1}\sum^{n}_{i=2}U_{1}\left(a^{t-1}_{1},a^{t-1}_{-1}[:i-1],a,a^{t-1}_{-1}[i+1:]\right)\right\}},\forall a\in\mathcal{A}.$